# Nonstandard Universes

### v1

@@ 1.1 Definition of a Universe
(Universe over $X$) If $X$ is a set, then $\mathbb{U}$ is said to be a universe over $X$ if and only if $X\in\mathbb{U}$, $\varnothing\not\in X$, and no element of a member of $X$ belongs to $\mathbb{U}$, i.e. $\bigcup X\cap\mathbb{U}=\varnothing$. Thus, from the point of view of $\mathbb{U}$, no member of $X$ has elements, yet none are the empty set — they are individuals.
@@ 1.2 Closure Properties of a Universe
    If $A\subseteq\mathbb{U}$ is finite, then $A\in\mathbb{U}$
@@ 1.3 Nonstandard Embeddings
Let $\mathcal{L}_{\in}$ denote a first-order language with equality $=$ and a single binary relation symbol $\in$. We assume that there is a countable collection of variables, and take as basic propositional connectives the connectives $\lnot$ (not) and $\land$ (and), and as basic quantifier $\forall$ (for all). The other connectives $\lor$ (or), $\to$ (then), $\leftrightarrow$ (if and only if) are defined in terms of ,$\land$,$\lnot$ , ¬ in the usual way, and the existential quantifier is defined in terms of ¬$\lnot$,$\forall$ , ∀ in the usual way.
@@
  1. (a)
    A transfer map for a set $X$ is a function ${}^{*}:\mathbb{U}\to\mathbb{V}$ between two universes $\mathbb{U},\mathbb{V}$ with the properties that:
@@
  2. (b)
    A transfer map $*:\mathbb{U}\to\mathbb{V}$ for $X$ is a nonstandard framework, or a nonstandard embedding, if there is a countable $C\in\mathbb{U}$ such that ${}^{\sigma}C$ is a proper subset of ${}^{*}C$, i.e. ${}^{\sigma}C\subsetneqq{}^{*}C$.
@@
Note that if * is a transfer map, then always ${}^{\sigma}A\subseteq{}^{*}A$, as $V(X)\vDash a\in A$ implies $V(Y)\vDash{}^{*}a\in{}^{*}A$. If $A$ is finite, then we shall see that ${}^{\sigma}A={}^{*}A$. If, however, the *-map is a nonstandard framework — so that $C$ is a proper subset of ${}^{*}C$ for some countable $C\in\mathbb{U}$— then it will transpire that ${}^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite. In that case, we can think of ${}^{*}A$ as a version of $A\in\mathbb{U}$ that lives in $\mathbb{V}$ — in that $A$ and ${}^{*}A$ satisfy the same bounded sentences — but where ${}^{*}A$ has additional elements that do not correspond to members of $A$.
Lemma 1.10
Suppose that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for $X$. Then $\mathbb{V}$ is a universe over ${}^{*}X$.
$\dashv$
@@
Thus without loss of generality, we may assume that, when working with a transfer map $*:\mathbb{U}\to\mathbb{V}$, the language $\mathcal{L}_{\in}$ is expanded to a language — denoted $\mathcal{L}_{\mathbb{U}}$ — which has a constant symbol $c_{a}$ for every entity $a\in\mathbb{U}$. Naturally, the constant symbol $c_{a}$ is to be interpreted as the entity $a$ in the model $\mathbb{U}$. If $*:\mathbb{U}\to\mathbb{V}$ is a transfer map, then $c_{a}$ will be interpreted as ${}^{*}a$ in $\mathbb{V}$. When we replace all occurrences of these constants in a formula $\varphi$ by their *-value, we obtain the *-transform ${}^{*}\varphi$ of the formula. Thus, for example
@@
In the next section, we will discuss some of the properties of nonstandard frameworks, should they exist. The question of existence is dealt with in Section 6.1, but, assuming familiarity with the model-theoretic concepts in Appendix A, this can be read now. We’ll have more to say about part (b) of Definition 1.9 later on. For now, note that this property is essential for nonstandard analysis to have any real power via the introduction of nonstandard objects.
@@ 2 Properties of the *-map
In this section, we assume that $*:\mathbb{U}\to\mathbb{V}$ is a transfer map between two universes $\mathbb{U},\mathbb{V}$.
@@
Observe that if $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$, then:
@@
      $\forall y\in\text{\rm ran}(R)\;(y\in R[C]\leftrightarrow\exists x\in C\;((x,y)\in R)$. if and only if
@@
If $*:\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$ and $V_{n}(X)\subseteq B\in\mathbb{U}$, then ${}^{*}V_{n}(X)=V_{n}({}^{*}X)\cap{}^{*}B$.
@@
The following result is crucial, as we wish to be able to talk about structures in our universes. For example, suppose that $*:V(\mathbb{R})\to V({}^{*}\mathbb{R})$ is a transfer map over the set $\mathbb{R}$ between two superstructures. When talking about the reals, we wish to take into account the operations and relations +,$+,\cdot,-,{}^{-1},\leq$ as well, so that we want to be able to talk about the model (,+$(\mathbb{R},+,\cdot,-,{}^{-1},\leq)$. This will transfer to a model (,+$({}^{*}\mathbb{R},{}^{*}+,{}^{*}\cdot,{}^{*}-,{}^{*}{}^{-1},{}^{*}\leq)$. The following result shows that these two models are elementarily equivalent, i.e. that they satisfy the same first-order sentences. Indeed, the *-map induces an elementary embedding from the one into the other:
@@
[…]
@@
Thus $*:\mathfrak{A}\to{}^{*}\mathfrak{A}$ is an elementary embedding.
@@ 3 Standard, Internal and External Objects
In this section, we suppose that we’re working with a transfer map $*:\mathbb{U}\to\mathbb{V}$ for a base set $X\in\mathbb{U}$.
@@ 3.1 Standard and Internal Objects
Here is the reason that internal objects play an important role: A transfer map $*:\mathbb{U}\to\mathbb{V}$ transfers the truth of bounded $\mathcal{L}_{\in}$-formulas from $\mathbb{U}$ to $\mathbb{V}$. Thus, for example if $a,b\in\mathbb{U}$, and $\varphi\equiv\forall x\in a\;\exists y\in b\;\exists z\in y\;\varphi(x,y,z,a,b)$, then with ${}^{*}\varphi\equiv\forall x\in{}^{*}a\;\exists y\in{}^{*}b\;\exists z\in y\;\varphi(x,y,z,{}^{*}a,{}^{*}b)$ we have $\mathbb{U}\vDash\varphi$ if and only if $\mathbb{V}\vDash{}^{*}\varphi$. Now to check if $\mathbb{V}\vDash{}^{*}\varphi$, we need merely check ${}^{*}\varphi$ over elements $x\in{}^{*}a,y\in{}^{*}b,z\in y$, i.e. we need only consider internal $x,y,z$. Then if ${}^{*}\varphi$ is true in $\mathbb{V}$, $\varphi$ is true in $\mathbb{U}$.
@@
But since $u\in{}^{*}\mathcal{P}(T)$, we have $u\subseteq{}^{*}T$, so that $u\cap{}^{*}T=u$. On the other hand, since $B\subseteq{}^{*}T$ (by transitivity of ${}^{*}T$), we have that $\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}\cap{}^{*}T=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$, and hence $u=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$. Finally, since $u\in{}^{*}\mathcal{P}(T)$, $u$ is internal.
@@
The map $*:(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ and the inclusion $({}^{*}\mathbb{U},\in)\hookrightarrow(\mathbb{V},\in)$ are bounded elementary embeddings.
@@
Now if $\varphi(x,y)$ is a bounded $\mathcal{L}_{\in}$-formula and $a,b\in\mathbb{U}$, then we have $\mathbb{U}\vDash\varphi(a,b)$ if and only if $\mathbb{V}\vDash\varphi({}^{*}a,{}^{*}b)$ if and only if ${}^{*}\mathbb{U}\vDash\varphi({}^{*}a,{}^{*}b)$, which shows that $*:(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ is a bounded elementary embedding.
@@ 3.2 Examples of Standard and Internal Objects
In this section, assume that $*:\mathbb{U}\hookrightarrow\mathbb{V}$ is a transfer map over $X$.
@@
$\bigcup\mathcal{A}=\{x\in{}^{*}T:\exists A\in\mathcal{A}\;(x\in A)\},$
which is internal by the internal definition principle. For $\bigcap\mathcal{A}\subseteq{}^{*}T$, just replace $\exists A\in\mathcal{A}\;(x\in A)$ by $\forall A\in\mathcal{A}\;(x\in A)$.
@@
$\text{\rm dom}(R)$ =$=\{x\in T:\exists y\in T\;(x,y)\in R\},$
$\text{\rm ran}(R)$ =$=\{y\in{}^{*}T:\exists x\in T\;(x,y)\in R\},$
$R^{-1}$ =$=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists v\in R\;(v=(x,y)\land w=(y,x))\},$
$R[C]$ =$=\{y\in{}^{*}T:\exists x\in C\;((c,y)\in R)\}.$
@@
$R\circ S=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists z\in{}^{*}T\;((x,y)\in S\land(y,z)\in R\land w=(x,z))\},$
where $(x,y)\in S$ is short for $\exists s\in S\;(s=(x,y))$, and $s=(x,y)$ is a bounded formula, by Lemma 2.1(c).
(d) $f$ is a binary relation, so $\text{\rm ran}(f)$ is internal. Since $f(a)\in\text{\rm ran}(f)$, $f(a)$ is internal, by transitivity of ${}^{*}\mathbb{U}$.
@@
Let $*:\mathbb{U}\hookrightarrow\mathbb{V}$ be a transfer map for $X$. Then
@@
Proof: Let $\mathcal{X}:=\{\mathcal{P}(A):A\in\mathcal{A}\}$, and let $T\in\mathbb{U}$ be transitive such that $\mathcal{X}\subseteq T$. Observe that
@@
Thus $P\in{}^{*}\mathcal{X}$ if and only if $P\in{}^{*}T$ and there is $A\in{}^{*}\mathcal{A}$ such that $P\cap{}^{*}T=\mathcal{P}(A)\cap{}^{*}T$. Since $P\in{}^{*}T$ and ${}^{*}T$ is transitive, we have $P=\mathcal{P}(A)\cap{}^{*}T$ for some $A\in{}^{*}\mathcal{A}$. It follows that $P\subseteq P_{A}$. Now if $B\in P_{A}$, then $B\subseteq A$ and there is $S\in\mathbb{U}$ such that $B\in{}^{*}S$. Since
@@
from which it follows that $B\in{}^{*}T$, and thus that $B\in T$. Hence $P=\mathcal{P}(A)\cap{}^{*}T$ is the set $P_{A}=\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ of all internal subsets of $A$. It therefore follows that $P\in{}^{*}\mathcal{X}$ if and only if $P=P_{A}$ for some $A\in{}^{*}\mathcal{A}$.
@@
$[\forall a\in\bigcup\mathcal{A}\;\exists A\in\mathcal{A}\;(a\in A)]\land[\forall A\in\mathcal{A}\;\forall a\in A\;(a\in\bigcup\mathcal{A})].$
@@
We often deal with unions, intersections and products of indexed families of sets. Suppose, for example, that $\mathcal{A}:=\{A_{i}:i\in I\}\in\mathbb{U}$ is an indexed family of sets in $\mathbb{U}$, and let $f:I\to\mathcal{A}:i\mapsto A_{i}$ be the indexing function, where $f\in\mathbb{U}$. Then ${}^{*}f:{}^{*}I\to{}^{*}\mathcal{A}$ is a function, with ${}^{*}f({}^{*}i)={}^{*}(f(i))={}^{*}A_{i}$ when $i\in I$. For $i\in{}^{*}I-I$, define ${}^{*}A_{i}:={}^{*}f(i)$ — but note that if $i\in{}^{*}I-I$, then ${}^{*}A_{i}$ is not necessarily the *-value of a member of $\mathbb{U}$. Now since $f$ is surjective, so is ${}^{*}f$, and hence ${}^{*}\mathcal{A}=\{{}^{*}A_{i}:i\in{}^{*}I\}$. Hence by the previous result,
@@
Proof: Consider the bounded formula $\varphi_{5}(f,A,B)$ of Lemma 2.1, which asserts that $f:A\to B$, i.e. that $f\in B^{A}$. Transfer of the bounded formula $\forall f\in B^{A}\;\varphi_{5}(f,A,B)$ shows that every member of ${}^{*}(B^{A})$ is an internal function ${}^{*}A\to{}^{*}B$. Conversely, suppose that $f\in{}^{*}B^{{}^{*}A}$ is internal. Let $C\in\mathbb{U}$ so that $f\in C$. Transfer of $\forall f\in C\;(\varphi_{5}(f,A,B)\to f\in B^{A})$ shows that $f\in{}^{*}(B^{A})$.
@@
$\mathbb{U}\vDash\forall P\in S\;\Big{(}\exists A\in\mathcal{A}\;\exists B\in\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in T)\Big{)},$
@@
$\mathbb{V}\vDash\forall P\in{}^{*}S\;\Big{(}\exists A\in{}^{*}\mathcal{A}\;\exists B\in{}^{*}\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in{}^{*}T)\Big{)}.$

### v2

@@ 1.1 Definition of a Universe
(Universe over $X$) If $X$ is a set, then $\mathbb{U}$ is said to be a universe over $X$ if and only if $\in\mathbb{U}, X\varnothing\not\in X$, and no element of a member of $X$ belongs to $\mathbb{U}$, i.e. $\bigcup X\cap\mathbb{U}=\varnothing$. Thus, from the point of view of $\mathbb{U}$, no member of $X$ has elements, yet none are the empty set — they are individuals.
@@ 1.2 Closure Properties of a Universe
    If $A\subseteq\mathbb{U}$ is finite, then $A\in\mathbb{U}$ .
@@ 1.3 Nonstandard Embeddings
Let $\mathcal{L}_{\in}$ denote a first-order language with equality == and a single binary relation symbol $\in$. We assume that there is a countable collection of variables, and take as basic propositional connectives the connectives $\lnot$ (not) and $\land$ (and), and as basic quantifier $\forall$ (for all). The other connectives $\lor$ (or), $\to$ (then), $\leftrightarrow$ (if and only if) are defined in terms of $\land,\lnot$ in the usual way, and the existential quantifier is defined in terms of $\lnot,\forall$ in the usual way.
In addition, $\exists!y\;\psi(y)$ abbreviates the formula $\exists y\;(\psi(y)\land\forall z\;(\psi(z)\to z=y))$, which states that $y$ is the unique element for which $\psi$ holds.
@@
A transfer map for a set $X$ is a function $\mathbb{U}\to\mathbb{V}$ between two universes $\mathbb{U},\mathbb{V}$ with the properties that:
@@
Definition 1.10
(Nonstandard Framework/Nonstandard Embedding)
A transfer map $\mathbb{U}\to\mathbb{V}$ for $X$ is a nonstandard framework, or a nonstandard embedding, if there is a countable $C\in\mathbb{U}$ such that $^{\sigma}C$ is a proper subset of $^{*}C$, i.e. $^{\sigma}C\subsetneqq{}^{*}C$.
$\Box$
Note that if * is a transfer map, then always $^{\sigma}A\subseteq{}^{*}A$, as $\mathbb{U}\vDash a\in A$ implies $\mathbb{V}\vDash{}^{*}a\in{}^{*}A$. If $A$ is finite, then we shall see that $^{\sigma}A={}^{*}A$. If, however, the *-map is a nonstandard framework — so that $C$ is a proper subset of $^{*}C$ for some countable $C\in\mathbb{U}$— then it will transpire that $^{\sigma}A\subsetneqq{}^{*}A$ whenever $A\in\mathbb{U}$ is infinite. In that case, we can think of $^{*}A$ as a version of $A\in\mathbb{U}$ that lives in $\mathbb{V}$ — in that $A$ and $^{*}A$ satisfy the same bounded sentences — but where $^{*}A$ has additional elements that do not correspond to members of $A$.
@@
Thus without loss of generality, we may assume that, when working with a transfer map $\mathbb{U}\to\mathbb{V}$, the language $\mathcal{L}_{\in}$ is expanded to a language — denoted $\mathcal{L}_{\mathbb{U}}$ — which has a constant symbol $c_{a}$ for every entity $a\in\mathbb{U}$. Naturally, the constant symbol $c_{a}$ is to be interpreted as the entity $a$ in the model $\mathbb{U}$. If $\mathbb{U}\to\mathbb{V}$ is a transfer map, then $c_{a}$ will be interpreted as $^{*}a$ in $\mathbb{V}$. When we replace all occurrences of these constants in a formula $\varphi$ by their *-value, we obtain the *-transform $^{*}\varphi$ of the formula. Thus, for example
@@
Lemma 1.12
Suppose that $\mathbb{U}\to\mathbb{V}$ is a transfer map for $X$. Then $\mathbb{V}$ is a universe over $^{*}X$.
$\dashv$
In the next section, we will discuss some of the properties of nonstandard frameworks, assuming that they exist. The question of existence is dealt with in Section 6.1, but, assuming familiarity with the model-theoretic concepts in Appendix A, this can be read now. We’ll have more to say about Definition 1.10 later on. For now, note that this property is essential for nonstandard analysis to have any real power via the introduction of nonstandard objects.
@@ 2 Properties of the *-map
In this section, we assume that $\mathbb{U}\to\mathbb{V}$ is a transfer map between two universes $\mathbb{U},\mathbb{V}$.
@@
Observe that if $\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$, then:
@@
      $\forall y\in\text{\rm ran}(R)\;(y\in R[C]\leftrightarrow\exists x\in C\;((x,y)\in R))$.
@@
If $\mathbb{U}\to\mathbb{V}$ is a transfer map for a set $X$ and $V_{n}(X)\subseteq B\in\mathbb{U}$, then $^{*}V_{n}(X)=V_{n}({}^{*}X)\cap{}^{*}B$.
@@
[…]
@@
Suppose that $\mathbb{U}\to\mathbb{V}$ is a transfer map for $X$, and that $A\in\mathbb{U}$. Let $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ be a model of a first-order language $\mathcal{L}$ (which need not be the language $\mathcal{L}_{\in}$ of $\mathbb{U}$). Then the restriction $\restriction_{A}$ of * to $A$ is an elementary embedding from $\mathfrak{A}$ into $^{*}\mathfrak{A}$.
@@
Thus $\mathfrak{A}\to{}^{*}\mathfrak{A}$ is an elementary embedding.
@@ 3 Standard, Internal and External Objects
In this section, we suppose that we’re working with a transfer map $\mathbb{U}\to\mathbb{V}$ for a base set $X\in\mathbb{U}$.
@@ 3.1 Standard and Internal Objects
Here is the reason that internal objects play an important role: A transfer map $\mathbb{U}\to\mathbb{V}$ transfers the truth of bounded $\mathcal{L}_{\in}$-formulas from $\mathbb{U}$ to $\mathbb{V}$. Thus, for example if $a,b\in\mathbb{U}$, and $\varphi\equiv\forall x\in a\;\exists y\in b\;\exists z\in y\;\psi(x,y,z,a,b)$, then with $^{*}\varphi\equiv\forall x\in{}^{*}a\;\exists y\in{}^{*}b\;\exists z\in y\;\psi(x,y,z,{}^{*}a,{}^{*}b)$ we have $\mathbb{U}\vDash\varphi$ if and only if $\mathbb{V}\vDash{}^{*}\varphi$. Now to check if $\mathbb{V}\vDash{}^{*}\varphi$, we need merely check $^{*}\varphi$ over elements $x\in{}^{*}a,y\in{}^{*}b,z\in y$, i.e. we need only consider internal $x,y,z$. Then if $^{*}\varphi$ is true in $\mathbb{V}$, $\varphi$ is true in $\mathbb{U}$.
@@
But since $u\in{}^{*}\mathcal{P}(T)$, we have $u\subseteq{}^{*}T$, so that $u\cap{}^{*}T=u$. On the other hand, since $B\subseteq{}^{*}T$ (by transitivity of $^{*}T$), we have that $\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}\cap{}^{*}T=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$, and hence $u=\{y\in B:\varphi(y,A_{1},\dots,A_{n})\}$. Finally, since $\in{}^{*}\mathcal{P}(T), uu$ is internal.
@@
The map $(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ and the inclusion $({}^{*}\mathbb{U},\in)\hookrightarrow(\mathbb{V},\in)$ are bounded elementary embeddings.
@@
Now if $\varphi(x,y)$ is a bounded $\mathcal{L}_{\in}$-formula and $a,b\in\mathbb{U}$, then we have $\mathbb{U}\vDash\varphi(a,b)$ if and only if $\mathbb{V}\vDash\varphi({}^{*}a,{}^{*}b)$ if and only if $^{*}\mathbb{U}\vDash\varphi({}^{*}a,{}^{*}b)$, which shows that $(\mathbb{U},\in)\hookrightarrow({}^{*}\mathbb{U},\in)$ is a bounded elementary embedding.
@@ 3.2 Examples of Standard and Internal Objects
In this section, assume that $\mathbb{U}\hookrightarrow\mathbb{V}$ is a transfer map over $X$.
@@
$\bigcup\mathcal{A}=\{x\in{}^{*}T:\exists y\in\mathcal{A}\;(x\in y)\},$
which is internal by the internal definition principle. For $\bigcap\mathcal{A}\subseteq{}^{*}T$, just replace $\exists y\in\mathcal{A}\;(x\in y)$ by $\forall y\in\mathcal{A}\;(x\in y)$.
@@
$\text{\rm dom}(R)$ $=\{x\in T:\exists y\in T\;((x,y)\in R)\},$
$\text{\rm ran}(R)$ $=\{y\in{}^{*}T:\exists x\in T\;((x,y)\in R)\},$
$R^{-1}$ $=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists v\in R\;(v=(x,y)\land w=(y,x))\},$
$R[C]$ $=\{y\in{}^{*}T:\exists x\in C\;((c,y)\in R)\}.$
Here, $(x,y)\in R$ is short for $\exists z\in R\;(z=(x,y))$, and $z=(x,y)$ is a bounded formula, by Lemma 2.1(c).
@@
$R\circ S=\{w\in{}^{*}(T\times T):\exists x\in{}^{*}T\;\exists y\in{}^{*}T\;\exists z\in{}^{*}T\;((x,y)\in S\land(y,z)\in R\land w=(x,z))\}.$
(f) $f$ is a binary relation, so $\text{\rm ran}(f)$ is internal. Since $\in\text{\rm ran}(f), f(a)f(a)$ is internal, by transitivity of $^{*}\mathbb{U}$.
@@
Let $\mathbb{U}\hookrightarrow\mathbb{V}$ be a transfer map for $X$. Then
@@
Proof: Let $\mathcal{X}:=\{\mathcal{P}(A):A\in\mathcal{A}\}$, and observe that $\mathcal{X}$ is the range of the map $\mathcal{A}\to\mathcal{P}\mathcal{P}(\bigcup\mathcal{A}):A\mapsto\mathcal{P}(A)$, so that $\mathcal{X}\in\mathbb{U}$. Let $T\in\mathbb{U}$ be transitive such that $\mathcal{X}\subseteq T$. Observe that
@@
Thus $P\in{}^{*}\mathcal{X}$ if and only if $P\in{}^{*}T$ and there is $A\in{}^{*}\mathcal{A}$ such that $P\cap{}^{*}T=\mathcal{P}(A)\cap{}^{*}T$. Since $P\in{}^{*}T$ and $^{*}T$ is transitive, we have $P=\mathcal{P}(A)\cap{}^{*}T$ for some $A\in{}^{*}\mathcal{A}$. It follows that $P\subseteq P_{A}$.
Now if $A\in{}^{*}\mathcal{A}$ and $B\in P_{A}$, then $B\subseteq A$ and there is $S\in\mathbb{U}$ such that $B\in{}^{*}S$. Since
@@
from which it follows that $B\in{}^{*}T$, and thus that $B\in P$. Hence $P=\mathcal{P}(A)\cap{}^{*}T$ is the set $P_{A}=\mathcal{P}(A)\cap{}^{*}\mathbb{U}$ of all internal subsets of $A$. It therefore follows that $P\in{}^{*}\mathcal{X}$ if and only if $P=P_{A}$ for some $A\in{}^{*}\mathcal{A}$.
@@
$\forall a\in\bigcup\mathcal{A}\;\exists A\in\mathcal{A}\;(a\in A)]\land[\forall A\in\mathcal{A}\;\forall a\in A\;(a\in\bigcup\mathcal{A})].$
@@
We often deal with unions, intersections and products of indexed families of sets. Suppose, for example, that $\mathcal{A}:=\{A_{i}:i\in I\}\in\mathbb{U}$ is an indexed family of sets in $\mathbb{U}$, and let $f:I\to\mathcal{A}:i\mapsto A_{i}$ be the indexing function, where $f\in\mathbb{U}$. Then $^{*}f:{}^{*}I\to{}^{*}\mathcal{A}$ is a function, with $^{*}f({}^{*}i)={}^{*}(f(i))={}^{*}A_{i}$ when $i\in I$. For $i\in{}^{*}I-I$, define $^{*}A_{i}:={}^{*}f(i)$but note that if $i\in{}^{*}I-I$, then $^{*}A_{i}$ is not necessarily the *-value of a member of $\mathbb{U}$. Now since $f$ is surjective, so is $^{*}f$, and hence $^{*}\mathcal{A}=\{{}^{*}A_{i}:i\in{}^{*}I\}$. Hence by the previous result,
@@
Proof: Consider the bounded formula $\varphi_{5}(f,A,B)$ of Lemma 2.1, which asserts that $f:A\to B$, i.e. that $f\in B^{A}$. Transfer of the bounded formula $\forall f\in B^{A}\;\varphi_{5}(f,A,B)$ shows that every member of $^{*}(B^{A})$ is an internal function $^{*}A\to{}^{*}B$. Conversely, suppose that $f\in{}^{*}B^{{}^{*}A}$ is internal. Let $C\in\mathbb{U}$ so that $f\in{}^{*}C$. Transfer of $\forall f\in C\;(\varphi_{5}(f,A,B)\to f\in B^{A})$ shows that $f\in{}^{*}(B^{A})$.
@@
$\mathbb{U}\vDash\forall P\in S\;\Big(\exists A\in\mathcal{A}\;\exists B\in\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in T\Big),$
@@
$\mathbb{V}\vDash\forall P\in{}^{*}S\;\Big(\exists A\in{}^{*}\mathcal{A}\;\exists B\in{}^{*}\mathcal{B}\;\varphi_{4}(P,A,B)\to P\in{}^{*}T\Big).$